# Graph Network Simulator MGN MeshGraphNet

### v1

@@
Denoising Diffusion Probalistic Model
@@
Hierarchical Contact Mesh Transforme
@@ Abstract
Graph-based learned simulators have emerged as a promising approach for simulating physical systems on unstructured meshes, offering speed and generalization across diverse geometries. However, they often struggle with capturing global phenomena, such as bending or long-range correlations, and suffer from error accumulation over long rollouts due to their reliance on local message passing and direct next-step prediction. We address these limitations by introducing the Rolling Diffusion-Batched Inference Network (ROBIN), a novel learned simulator that integrates two key innovations: (i) Rolling Diffusion, a parallelized inference scheme that amortizes the cost of diffusion-based refinement across physical time steps by overlapping denoising steps across a temporal window. (ii) A Hierarchical Graph Neural Network built on algebraic multigrid coarsening, enabling multiscale message passing across different mesh resolutions. This architecture, implemented via Algebraic-hierarchical Message Passing Networks, captures both fine-scale local dynamics and global structural effects critical for phenomena like beam bending or multi-body contact. We validate ROBIN on challenging $2$D and $3$D solid mechanics benchmarks involving geometric, material, and contact nonlinearities. ROBIN achieves state-of-the-art accuracy on all tasks, substantially outperforming existing next-step learned simulators while reducing inference time by up to an order of magnitude compared to standard diffusion simulators.
@@ 1 Introduction
Physical simulations enable many engineering and scientific fields to gain quick insights into complex systems or to evaluate design decisions. Conventional simulations model the physical system using Partial Differential Equations. Usually, the PDE is discretized by numerical methods, such as the  Finite Element Method (FEM[1], the  Finite Volume Method (FVM[2] or the  Finite Difference Method (FDM[3]. This process reduces the need for cumbersome, resource-intensive real-world experiments. Recent research aims to speed up simulation with  Machine Learning (ML)-based models [4, 5]. These learned simulators promise to allow researchers and practitioners to evaluate large amounts of virtual, simulated experiments. These simulations in turn unlock applications in engineering design and manufacturing optimization [6, 7, 8].
@@
This work aims to improve learned simulators, focusing on simulations from nonlinear solid mechanics as a representative class of examples. We combine recent image-based Denoising Diffusion Probalistic Models [9, 10, 11] with  Hierarchical Graph Neural Network (HGNN[12, 13, 14]. While diffusion has shown promising results on images [15, 16, 17], audio [18, 19] and even policy learning for robotics [20, 21], it suffers from cost-intensive inference due to its iterative denoising procedure. We alleviate this high inference cost on time-dependent domains with   Rolling Diffusion-Batched Inference (ROBI), a novel scheduling scheme that batches denoising steps of consecutive time steps. ROBI already starts denoising future prediction steps by using partially refined previous steps. This process reduces the number of model evaluations to the number of time steps and preserves the time-shift equivarinace of Markovian systems. ROBI only affects the inference processes, allowing us to utilize conventional, parallelized DDPM training.
@@
Figure 1: Overview of ROBINROBIN coarses the fine mesh multiple times with AMG to create a graph hierarchy. ROBIN predicts the denoising velocity batch $\mathbf{v}_{\theta,k_{\text{B}}}^{t_{\text{B}}}$ using AMPNs, given a batch of noisy samples $\Delta{\mathbf{u}}^{t_{\text{B}}}_{k_{\text{B}}}$ and clean samples $\Delta\tilde{\mathbf{u}}^{t_{\text{B}}}_{0|k_{\text{B}}+1}$ as input. The prediction $\mathbf{v}_{\theta,k_{\text{B}}}^{t_{\text{B}}}$ is used for a denoising step of the noisy sample $\Delta{\mathbf{u}}^{t_{\text{B}}}_{k_{\text{B}}-1}$ and a state update $\Delta\tilde{\mathbf{u}}^{t_{\text{B}}}_{0|k_{\text{B}}}$.
@@
We combine ROBI with a diffusion HGNN to form the  Rolling Diffusion-Batched Inference Network (ROBIN), which significantly accelerates simulation while improving predictive accuracy. ROBIN constitutes the first diffusion-based refiner for simulating physical dynamics on unstructured meshes, surpassing the current accuracy ceiling of hierarchical Graph Neural Networks. We train ROBINs on three challenging $2$D and $3$D solid mechanics datasets involving geometric, material, and contact nonlinearities. Across all datasets, ROBIN significantly improves over state-of the-art mesh-based simulators [5, 22] in terms of predictive accuracy. We further find that our proposed inference method, ROBI, speeds up diffusion-based inference for learned simulations by up to an order of magnitude while maintaining accuracy. [^1] Code and datasets will be open-sourced upon acceptance.
@@
To summarize, our contributions we i) propose ROBI, a novel inference scheduling scheme for diffusion-based simulators that amortizes denoising across time steps, reducing inference to a single model evaluation per step while preserving time-shift equivariance; ii) introduce ROBIN, a diffusion-based HGNN for nonlinear solid mechanics that combines multiscale message passing with ROBI to provide fast, accurate diffusion-based simulations; iii) demonstrate state-of-the-art performance on challenging $2$D and $3$D benchmarks, outperforming existing simulators in both accuracy and runtime.
@@ 2 Related Work
Simulating Complex Physics. Simulating complex physical systems often requires numerical solvers, such as the FEM [23, 24, 1], the FVM [2], or the FDM [3]. While accurate, numerical solvers scale poorly with problem complexity, often requiring multiple hours of even days for a single rollout on a modern workstation. Recent work shows that ML-based models are able to learn such numerical simulations from data [25, 26, 4, 5, 27, 9, 22, 10]. ML-based models provide speed-ups of one to two orders of magnitude while being fully differentiable, which accelerates downstream applications such as design [6] or manufacturing process optimization [7, 8]. Many learned simulators operate autoregressively. They mimic numerical solvers by using their own predictions to estimate the residual between successive time steps [4, 5, 28]. Similarly, Neural ODEs [29, 30] predict time derivatives and advance solutions via numerical integration. In contrast to these supervised approaches, Physics-Informed Neural Networks [31] directly operate on a PDEs loss function to train Multilayer Perceptrons [32, 33, 8]Convolutional Neural Networks [34, 35] or GNN [36, 37]. Finally, Neural Operators aim to learn mesh-independent solution operators [38, 39, 40, 27, 41]. ROBIN is an autoregressive learned simulator that replaces direct next-step prediction with multiple denoising diffusion steps, leveraging generative inference to improve prediction accuracy.
Learning Mesh-based Simulations with Graph Neural Network (GNN). Pfaff et al. [5] introduced MeshGraphNets, a Message Passing Network (MPN)-based simulator that encodes simulation states as graphs using mesh connectivity and physical proximity. While accurate for small problems, MeshGraphNet (MGN) does not scale well, as its receptive field is limited by the number of message-passing steps in the MPN. To address this issue, recent work expands the receptive field via global attention [42, 43, 44] or hierarchical mesh representations using Graph Convolutional Networks [45] and MPNs [46, 13]. Extensions further improve efficiency and accuracy through rotation equivariance [14], hierarchical edge design [47], bi-stride pooling [48], and attention mechanisms acting on edges and across hierarchies [22]. [49] leverages  Adaptive Mesh Refinement (AMR) to create mesh hierarchies for multi-scale GNNs. Existing methods are generally trained using a next-step Mean Squared Error (MSE) loss, which favors learning lower solutions frequencies at the cost of accuracy in higher frequency bands that have less impact on the loss [9]. However, autoregressive models trained with a MSE loss often overlook low-amplitude frequencies [9]. Our approach is orthogonal, coupling hierarchical GNN with denoising diffusion models. This approach allows to take advantage of the large receptive field of multi-scale GNNs while pushing accuracy toward diffusion limits.
@@
While next-step simulators trained with MSE loss capture high-amplitude, low-frequency components, they often miss low-amplitude components [9]. PDE-Refiners (PDE-Refiners) addresses this via iterative refinement, improving long-horizon accuracy of grid-based CNN simulators. We extend this idea to unstructured meshes by combining algebraic mesh coarsening [53, 54] with hierarchical GNNs [52, 22] with shared layers. We further introduce a time-parallel denoising scheme at inference, removing the speed bottleneck while maintaining accuracy. Unlike diffusion-based CNN simulators that require $K$ model evaluations per physical step $T$ [11, 55, 9], our method requires only a single hierarchical GNN call per step post warm-up, reducing inference from $\mathcal{O}(KT)$ to $\mathcal{O}(T)$. Compared to video-based approaches [10, 56], which need to jointly process $N$ steps and learn a time-dependent denoiser with high memory cost, our model, ROBIN, leverages time-translation invariance to train a time-independent denoiser with only one time step in memory. As such, ROBIN can be applied autoregressively and can freely interpolate between fully parallel denoising and memory-efficient one-step denoising at inference time. It also predicts state residuals instead of states, significantly improving long-horizon rollout fidelity.
@@ 3   Rolling Diffusion-Batched Inference Network (ROBIN)
Graph Network Simulators (GNSs) for Mesh-based Simulations. We consider solving PDEs for physical quantities $\mathbf{u}(\mathbf{x},t)$ that change over time $t\in[0,T]$ and inside a time-dependent domain $\mathbf{x}(t)\in\Omega(t)$. We focus on simulations on meshes, where $\mathcal{G}=(\mathcal{V},\mathcal{E}^{\text{M}})$ denotes the mesh graph and the graph nodes $\mathcal{V}$ and the graph edges $\mathcal{E}^{\text{M}}$ correspond to mesh nodes and mesh edges. We seek solutions $\mathbf{u}_{i}(t)=\mathbf{u}(\mathbf{x}_{i},t)$ at discrete node locations $\mathbf{x}_{i}(t)\in\Omega(t)$. To obtain discretized PDEs, usually numerical methods, such as the FEM, are applied that define the discretization of spatial operators, such as gradients $\partial\mathbf{u}(\mathbf{x},t)/\partial\mathbf{x}$. Given the discretized operator $\mathcal{F}$, the PDE simplifies to an time-dependent Ordinary Differential Equation and requires solving $\partial\mathbf{u}_{i}/\partial t=\mathcal{F}(t,\mathbf{x}_{i},\mathbf{u}_{i})~{}\text{.}$ We can solve such systems using numerical time discretization schemes. In this work we use a simple Euler forward discretization $\mathbf{u}_{i}^{t{+}1}=\mathbf{u}_{i}^{t}+\Delta t~{}\mathcal{F}(t,\mathbf{x}_{i}^{t},\mathbf{u}_{i}^{t})~{}\text{,}$ and set $\Delta t=1$. We extend  PDE-Refiner (PDE-Refiner[9] to Lagrangian systems, where the domain $\Omega=\Omega^{t}$ and node locations $\mathbf{x}_{i}=\mathbf{x}_{i}^{t}\in\Omega^{t}$ evolve over time. Here, we predict the solution $\mathbf{u}_{i}^{t}$ at time step $t$ by learning to reverse a probabilistic diffusion process[57] conditioned on the solution history. The proposed methods also apply without any restriction to Eulerian systems, where the domains are time-independent.
3.1 Denoising Diffusion Probalistic Models (DDPMs) for time-dependent simulations
@@
$p_{\theta}(\mathbf{u}^{t}_{k-1}|\mathbf{u}^{t}_{k}):=\mathcal{N}(\mathbf{u}^{t}_{k-1};\mu_{\theta}(\mathbf{u}^{t}_{k},k),\Sigma_{\theta}(\mathbf{u}^{t}_{k},k))~{}\text{,}$
where the mean $\mu_{\theta}$ and covariance $\Sigma_{\theta}$ are predicted by a learned model with parameters $\theta$. The covariance is assumed to be isotropic and given as $\sigma_{k}^{2}\mathbf{I}=(\frac{1-\bar{\alpha}_{k-1}}{1-\bar{\alpha}_{k}}\beta_{k})\mathbf{I}$ [57] with $\alpha_{i}=1-\beta_{i}$ and $\bar{\alpha}_{k}=\prod_{i=1}^{k}\alpha_{i}$. The reverse denoising process consists of $K$ iterative diffusion steps and starts from $k=K$. To facilitate faster denoising, we follow the DDPM formulation of [9] and use an exponential $\beta_{k}$ scheduler. We train the model to predict the denoising velocity, i.e., the v-prediction target
@@
given gaussian noise $\mathbf{\epsilon}^{t}$ [58]. We predict the solution autoregressevily, conditioning the model on the last time step solution $\mathbf{u}^{t-1}_{0}$ and additional conditioning features $\mathbf{c}^{t-1}$ to predict the current denoising velocity $\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k)=\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k,\mathbf{u}^{t-1}_{0})$. We define the target as $\mathbb{E}_{\mathbf{u}^{t}_{0},\mathbf{\epsilon}^{t},k}\left[\left\|\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k,\mathbf{u}^{t-1}_{0})-\mathbf{v}_{k}^{t}\right\|^{2}\right]$, which corresponds to a MSE loss over predicted denoising velocities [9]. The first denoising step is defined such that $\bar{\alpha}_{K}\approx 0$, which simplifies the v-prediction target to $\mathbf{v}_{k}^{t}\approx-\mathbf{u}^{t}_{0}$ (cf. Equation 1). The noisy sample $\mathbf{u}^{t}_{k}\approx\mathbf{\epsilon}^{t}$ corresponds to Gaussian Noise and is uninformative. Hence, for $k=K$ the model target converges to $\left\|\mathbf{v}_{\theta}(\mathbf{u}^{t}_{K},K,\mathbf{u}^{t-1}_{0}))-\mathbf{v}_{k}^{t}\right\|^{2}\approx\left\|\mathbf{v}_{\theta}(\mathbf{\epsilon}^{t},K,\mathbf{u}^{t-1}_{0}))+\mathbf{u}^{t}_{0}\right\|^{2}$, i.e., the MSE objective of one step models. Consequently, the predicted solution after the first denoising step $k=K$ mirrors that of one-step models [9]. More importantly, due to the exponential noise scheduler, each denoising step focuses on different amplitude and frequency levels of the solution [9], with later denoising steps increasingly paying attention to higher frequencies.
Rolling Diffusion-Batched Inference (ROBI). Conventional diffusion inference requires $K$ model calls, each corresponding to a denoising step, per simulation time step [9]. Figure 2 a) shows an example. Thus, inference is roughly $K$ times slower than one-step models [5]. We propose  Rolling Diffusion-Batched Inference (ROBI) to accelerate inference in DDPM-based autoregressive simulators. Given a velocity prediction $\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k,\mathbf{u}^{t-1}_{0})$ at the denoising step $k$, we reconstruct a partially refined prediction $\tilde{\mathbf{u}}^{t}_{0|k}=\sqrt{\bar{\alpha}_{k}}\mathbf{u}^{t}_{k}-\sqrt{1-\bar{\alpha}_{k}}\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k,\mathbf{u}^{t-1}_{0})$  [58].
@@
Figure 2: Overview of a) conventional autoregressive diffusion inference and b) ROBI. a) Conventional inference denoises the entire state of a physical time step at once before it shifts to the next time step (see One Step Denosing). The Iterative Inference requires $K$ model calls per time step, where $K$ denotes the number of diffusion steps. b) ROBI parallelizes denoising steps across physical time, processing up to $K$ time steps batched, and reconstruct the physical states with the clean sample prediction subsequently (see Batched Denoising Step). This process allows Rolling-Batched Inference after the initial warm-up, reducing the number of model calls to one per time step.

### v2

@@
Denoising Diffusion Probabilistic Model
@@
Hierarchical Contact Mesh Transformer
@@
BSMS
Bi-Stride Multi-Scale GNN
10 diffusion steps
10 diffusion steps
5 diffusion steps
5 diffusion steps
w/o shared layer
w/o shared layer
w/o diffusion
w/o diffusion
w/o hierarchy
w/o hierarchy
State prediction
State prediction
HCMT model
HCMT model
@@ Abstract
Graph-based learned simulators have emerged as a promising approach for simulating physical systems on unstructured meshes, offering speed and generalization across diverse geometries. However, they often struggle with capturing global phenomena, such as bending or long-range correlations usually occurring in solid mechanics, and suffer from error accumulation over long rollouts due to their reliance on local message passing and direct next-step prediction. We address these limitations by introducing the Rolling Diffusion-Batched Inference Network (ROBIN), a novel learned simulator that integrates two key innovations: (i) Rolling Diffusion-Batched Inference (ROBI), a parallelized inference scheme that amortizes the cost of diffusion-based refinement across physical time steps by overlapping denoising steps across a temporal window. (ii) A Hierarchical Graph Neural Network built on algebraic multigrid coarsening, enabling multiscale message passing across different mesh resolutions. This architecture, implemented via Algebraic-hierarchical Message Passing Networks, captures both fine-scale local dynamics and global structural effects critical for phenomena like beam bending or multi-body contact. We validate ROBIN on challenging 22D and 33D solid mechanics benchmarks involving geometric, material, and contact nonlinearities. ROBIN achieves state-of-the-art accuracy on all tasks, substantially outperforming existing next-step learned simulators while reducing inference time by up to an order of magnitude compared to standard diffusion simulators.
@@ 1 Introduction
Physical simulations enable many engineering and scientific fields to gain quick insights into complex systems or to evaluate design decisions. Conventional simulations model the physical system using Partial Differential Equations. Usually, the PDE is discretized by numerical methods, such as the  Finite Element Method (FEM[1], the  Finite Volume Method (FVM[2], or the  Finite Difference Method (FDM[3]. This process reduces the need for cumbersome, resource-intensive real-world experiments. Recent research aims to speed up simulation with  Machine Learning (ML)-based models [4, 5]. These learned simulators promise to allow researchers and practitioners to evaluate large amounts of virtual, simulated experiments. These simulations in turn unlock applications in engineering design and manufacturing optimization [6, 7, 8].
@@
This work aims to improve learned simulators, focusing on simulations of nonlinear solid mechanics as a representative class of examples. We combine recent image-based Denoising Diffusion Probabilistic Models [9, 10, 11] with Hierarchical Graph Neural Networks [12, 13, 14] (cf. Figure˜1). While diffusion has shown promising results on images [15, 16, 17], audio [18, 19] and even policy learning for robotics [20, 21], it suffers from cost-intensive inference due to its iterative denoising procedure. We alleviate this high inference cost on time-dependent domains with  Rolling Diffusion-Batched Inference (ROBI), a novel scheduling scheme that batches denoising steps of consecutive time steps. ROBI already starts denoising future prediction steps by using partially refined previous steps. This process reduces the number of model evaluations to the number of time steps and preserves the time-shift equivariance of Markovian systems. ROBI only affects the inference process, allowing us to utilize conventional, parallelized DDPM training.
@@
Figure 1: Overview of a ROBIN prediction. ROBIN coarsens the fine mesh multiple times with AMG to create a graph hierarchy. ROBIN predicts the denoising velocity $\mathbf{v}_{\theta}(\Delta\mathbf{u}^{t}_{k},k,\mathbf{u}^{t-1}_{0})$ at time step $t$ using AMPNs, given a noisy residual sample $\Delta{\mathbf{u}}^{t}_{k}$, the diffusion step $k$, and a previous state $\mathbf{u}^{t-1}_{0}$. The prediction is used to draw a new noisy residual sample $\Delta{\mathbf{u}}^{t}_{k-1}$ and to update the state $\mathbf{u}^{t}_{0}$.
@@
We combine ROBI with DDPMs and  Algebraic-hierarchical Message Passing Networks (AMPNs) to form the  Rolling Diffusion-Batched Inference Network (ROBIN), which significantly accelerates simulation while improving predictive accuracy. ROBIN constitutes the first diffusion-based refiner for simulating physical dynamics on unstructured meshes, surpassing the current accuracy ceiling of HGNNs. We train ROBINs on three challenging 22D and 33D solid mechanics datasets involving geometric, material, and contact nonlinearities. Across all datasets, ROBIN significantly improves over state-of-the-art mesh-based simulators [5, 22] in terms of predictive accuracy. Leveraging DDPMsROBIN accurately captures low-frequency global solution modes while resolving high-frequency components. We further find that our proposed inference method, ROBI, speeds up diffusion-based inference for learned simulations by up to an order of magnitude while maintaining accuracy. In addition, the AMPN architecture of ROBIN enables efficient transfer to much larger meshes, while maintaining near-FEM accuracy. [^1] Project page, code and datasets are available at https://tbswrth.github.io/ROBIN.
@@
To summarize, we i) propose ROBI, a novel inference scheduling scheme for diffusion-based simulators that amortizes denoising across time steps, reducing inference to a single model evaluation per step while preserving time-shift equivariance; ii) introduce ROBIN, a diffusion-based HGNN for nonlinear solid mechanics that combines multiscale message passing with ROBI to provide fast, accurate diffusion-based simulations; iii) demonstrate state-of-the-art performance on challenging 22D and 33D benchmarks, outperforming existing simulators in both accuracy and runtime.
@@ 2 Related Work
Simulating Complex Physics. Simulating complex physical systems often requires numerical solvers, such as the FEM [23, 24, 1], the FVM [2], or the FDM [3]. While accurate, numerical solvers scale poorly with problem complexity, often requiring multiple hours or even days for a single rollout on a modern workstation. Recent work shows that ML-based models are able to learn such numerical simulations from data [25, 26, 4, 5, 27, 9, 22, 10]. ML-based models provide speed-ups of one to two orders of magnitude while being fully differentiable, which accelerates downstream applications such as design [6] or manufacturing process optimization [7, 8]. Many learned simulators operate autoregressively. They mimic numerical solvers by using their own predictions to estimate the residual between successive time steps [4, 5, 28]. Similarly, Neural ODEs [29, 30] predict time derivatives and advance solutions via numerical integration. In contrast to these supervised approaches, Physics-Informed Neural Networks [31] directly operate on a PDE loss function to train Multilayer Perceptrons [32, 33, 8]Convolutional Neural Networks [34, 35] or Graph Neural Networks [36, 37]. Finally, Neural Operators aim to learn mesh-independent solution operators [38, 39, 40, 27, 41]. ROBIN is an autoregressive learned simulator that replaces direct next-step prediction with multiple denoising diffusion steps, leveraging generative inference to improve prediction accuracy.
Learning Mesh-based Simulations with Graph Neural Network (GNN). Pfaff et al. [5] introduced MeshGraphNets, a Message Passing Network (MPN)-based simulator that encodes simulation states as graphs using mesh connectivity and physical proximity. While accurate for small problems, MeshGraphNets does not scale well, as its receptive field is limited by the number of message-passing steps in the MPN. To address this issue, recent work expands the receptive field via global attention [42, 43, 44] or hierarchical mesh representations using Graph Convolutional Networks [45] and MPNs [46, 13]. Extensions further improve efficiency and accuracy through rotation equivariance [14], hierarchical edge design [47], bi-stride pooling [48], and attention mechanisms acting on edges and across hierarchies [22]. [49] leverages  Adaptive Mesh Refinement (AMR) to create mesh hierarchies for multi-scale GNNs. Complementary, Physics-Informed MeshGraphNets [37] integrate FEM-based training to improve accuracy and robustness. Existing methods are generally trained using a next-step Mean Squared Error (MSE) loss, which favors learning lower solution frequencies at the cost of accuracy in higher frequency bands that have less impact on the loss [9]. However, autoregressive models trained with a MSE loss often overlook low-amplitude frequencies [9]. Our approach is orthogonal, coupling hierarchical GNN with denoising diffusion models. This approach takes advantage of the large receptive field of multi-scale GNNs while pushing accuracy toward diffusion limits.
@@
While next-step simulators trained with MSE loss capture high-amplitude, low-frequency components, they often miss low-amplitude components [9]. PDERefs address this via iterative refinement, improving long-horizon accuracy of grid-based CNN simulators. We extend this idea to unstructured meshes by combining algebraic mesh coarsening [53, 54] with hierarchical GNNs [52, 22] with shared layers. We further introduce a time-parallel denoising scheme at inference, removing the speed bottleneck while maintaining accuracy. Unlike diffusion-based CNN simulators that require $K$ model evaluations per physical step [11, 55, 9], our method requires only a single hierarchical GNN call per step post warm-up, reducing inference costs over $T$ time steps from $\mathcal{O}(KT)$ to $\mathcal{O}(T)$. Compared to video-based approaches [10, 56], which need to jointly process $N$ steps and learn a time-dependent denoiser with high memory cost, our model, ROBIN, leverages time-translation invariance to train a time-independent denoiser with only one time step in memory. As such, ROBIN can be applied autoregressively and can freely interpolate between fully parallel denoising and memory-efficient one-step denoising at inference time. It also predicts state residuals instead of states, significantly improving long-horizon rollout fidelity.
@@ 3  Rolling Diffusion-Batched Inference Network (ROBIN)
Graph Network Simulators (GNSs) for Mesh-based Simulations. We consider solving PDEs for physical quantities $\mathbf{u}(\mathbf{x},t)$ that change over time $t\in[0,T]$ and inside a time-dependent domain $\mathbf{x}(t)\in\Omega(t)$. We focus on simulations on meshes, where $\mathcal{G}=(\mathcal{V},\mathcal{E}^{\text{M}})$ denotes the mesh graph and the graph nodes $\mathcal{V}$ and the graph edges $\mathcal{E}^{\text{M}}$ correspond to mesh nodes and mesh edges. We seek solutions $\mathbf{u}_{i}(t)=\mathbf{u}(\mathbf{x}_{i},t)$ at discrete node locations $\mathbf{x}_{i}(t)\in\Omega(t)$. To obtain discretized PDEs, usually numerical methods, such as the FEM, are applied that define the discretization of spatial operators, such as gradients $\partial\mathbf{u}(\mathbf{x},t)/\partial\mathbf{x}$. Given the discretized operator $\mathcal{F}$, the PDE simplifies to a time-dependent Ordinary Differential Equation and requires solving $\partial\mathbf{u}_{i}/\partial t=\mathcal{F}(t,\mathbf{x}_{i},\mathbf{u}_{i})~\text{.}$ We can solve such systems using numerical time discretization schemes. In this work we use a simple Euler forward discretization $\mathbf{u}_{i}^{t{+}1}=\mathbf{u}_{i}^{t}+\Delta t~\mathcal{F}(t,\mathbf{x}_{i}^{t},\mathbf{u}_{i}^{t})~\text{,}$ and set $\Delta t=1$. We extend PDE-Refiner [9] to Lagrangian systems, where the domain $\Omega^{t}$ and node locations $\mathbf{x}_{i}^{t}\in\Omega^{t}$ evolve over time. Here, we predict the solution $\mathbf{u}_{i}^{t}$ at time step $t$ by learning to reverse a probabilistic diffusion process [57] conditioned on the previous state of time step t1t-1. The proposed methods also apply without any restriction to Eulerian systems, where the domains are time-independent.
3.1 Denoising Diffusion Probabilistic Models (DDPMs) for time-dependent simulations
@@
$p_{\theta}(\mathbf{u}^{t}_{k-1}|\mathbf{u}^{t}_{k}):=\mathcal{N}(\mathbf{u}^{t}_{k-1};\mu_{\theta}(\mathbf{u}^{t}_{k},k),\Sigma(\mathbf{u}^{t}_{k},k))~\text{,}$
consisting of $K$ iterative diffusion steps and starting from $k=K$. The mean $\mu_{\theta}$ depends on the prediction of a learned model with parameters $\theta$. The covariance $\Sigma$ is assumed to be isotropic and given as $\Sigma=\sigma_{k}^{2}\mathbf{I}=(\frac{1-\bar{\alpha}_{k-1}}{1-\bar{\alpha}_{k}}\beta_{k})\mathbf{I}$ [57] with $\alpha_{i}=1-\beta_{i}$ and $\bar{\alpha}_{k}=\prod_{i=1}^{k}\alpha_{i}$. We train the model to predict the denoising velocity, i.e., the v-prediction target
@@
given gaussian noise $\mathbf{\epsilon}^{t}$ [58]. Since this target smoothly interpolates between $\mathbf{u}_{0}^{t}$ as $\bar{\alpha}_{k}\approx 0$ and $\boldsymbol{\epsilon}^{t}$ ($\bar{\alpha}_{k}\approx 1$), it emphasizes predicting the clean sample in early (high-noise) steps and the noise in later (high-signal) steps, simplifying learning. The model predicts the denoising velocity $\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k)=\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k,\mathbf{u}^{t-1}_{0})$ autoregressively, conditioned the model on the last time step solution $\mathbf{u}^{t-1}_{0}$. The training objective is then defined as $\mathbb{E}_{\mathbf{u}^{t}_{0},\mathbf{\epsilon}^{t},k}\left[\left\|\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k,\mathbf{u}^{t-1}_{0})-\mathbf{v}_{k}^{t}\right\|^{2}\right]$ [9]. To facilitate faster denoising, we follow the DDPM formulation of [9] and use an exponential $\beta_{k}$ scheduler.
@@
Figure 2: Overview of a) conventional autoregressive diffusion inference and b) ROBI. a) Conventional inference denoises the entire state of a physical time step at once before it shifts to the next time step (see One Step Denoising). The Iterative Inference requires $K$ model calls per time step, where $K$ denotes the number of diffusion steps. b) ROBI parallelizes denoising steps across physical time, processing up to $K$ time steps batched, and reconstruct the physical states with the clean sample prediction subsequently (see Batched Denoising Step). This process allows Rolling-Batched Inference after the initial warm-up, reducing the number of model calls to one per time step.
The first denoising step is defined such that $\bar{\alpha}_{K}\approx 0$, which simplifies the v-prediction target to $\mathbf{v}_{k}^{t}\approx-\mathbf{u}^{t}_{0}$ (cf. Equation˜1). The noisy sample $\mathbf{u}^{t}_{k}\approx\mathbf{\epsilon}^{t}$ corresponds to Gaussian Noise and is uninformative. Hence, for $k=K$ the model target converges to $\left\|\mathbf{v}_{\theta}(\mathbf{u}^{t}_{K},K,\mathbf{u}^{t-1}_{0}))-\mathbf{v}_{k}^{t}\right\|^{2}\approx\left\|\mathbf{v}_{\theta}(\mathbf{\epsilon}^{t},K,\mathbf{u}^{t-1}_{0}))+\mathbf{u}^{t}_{0}\right\|^{2}$. This MSE objective of the first denoising step $k=K$ mirrors the training objective of one-step models [9], i.e., auto-regressive models that predict the solution of the next time directly with a single prediction step. Hence, we expect similar accuracy as one-step models during the first diffusion steps. However, the DDPM-based model refines the initial prediction iteratively at each time step, improving the accuracy of the solution further. Note that, due to the noise scheduler, each denoising step focuses on different amplitude and frequency levels of the solution [9], with later denoising steps increasingly paying attention to higher frequencies.
Rolling Diffusion-Batched Inference (ROBI). Conventional diffusion inference requires $K$ model calls, each corresponding to a denoising step, per simulation time step [9]. Figure˜2 a) shows an example. Thus, inference is roughly $K$ times slower than one-step models [5]. We propose  Rolling Diffusion-Batched Inference (ROBI) to accelerate inference in DDPM-based autoregressive simulators. Given a velocity prediction $\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k,\mathbf{u}^{t-1}_{0})$ at the denoising step $k$, we reconstruct a partially refined prediction $\tilde{\mathbf{u}}^{t}_{0|k}=\sqrt{\bar{\alpha}_{k}}\mathbf{u}^{t}_{k}-\sqrt{1-\bar{\alpha}_{k}}\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k,\mathbf{u}^{t-1}_{0})$, following [58].